1. Three cards are randomly selected from a standard 52 card deck without replacement. Find
the probabilities of the following events:
(a) (3 points) All cards are Aces.
(b) (3 points) All cards are Diamonds.
(c) (3 points) First is a King, the second is an Ace and the third one is another Ace.
2. A number is selected randomly from the list f1; 2; 3; 4; 5; 6; 7; 8; 9; 10g. Consider the following
events:
A= The number selected is less or equal to 7. B= The number selected is more than 2.
(a) (3 points) Compute P(A) and P(B).
(b) (3 points) Explain the meaning in this situation of the event A \ B. Find P(A \ B).
(c) (3 points) Explain the meaning in this situation of the event A [ B. Find P(A [ B).
4. A baseball player hits the ball 40% of the times. What is the probability of getting:
(a) (3 points) exactly four hits in 9 opportunities.
(b) (3 points) at most four hits in 9 opportunities.
(c) (5 points) If the random variable X represents the amount of hits in 9 opportunities, what
kind of distribution of probability is represented in this experiment? what is the expected
value  and standard deviation ?

Answer by edjones(8007)

1.
a)
4C3/52C3
=4/22100
=1/5525
.
b)
13C3/52C3
=286/22100
=11/850
.
c)
4/52 * 4/51 * 3/50
=2/5525
.
Ed